An electric heater is used to heat 0.5kg of water in a kettle of heat capacity 400j/k, kept at 20degrees celsius.

what mass of water is boiled away in 600s? (Neglect heat losses)


1
Expert's answer
2021-11-21T17:32:18-0500

The heat balance equation gives

"Q=cm\Delta T+L\Delta m=P\tau"

The mass of water that is boiled away

"\Delta m=\frac{P\tau-cm\Delta T}{L}"

"\Delta m=\frac{2000*600-4200*0.5*80}{2.26*10^6}=0.46\:\rm kg"